OpenStudy (vortish):

find the slope intercept form for the line satisfying the following conditions passing thru (7,-6)(6,-5) the slope intercept form for the line is

OpenStudy (anonymous):

Step 1. Find the slope between these two points.

OpenStudy (anonymous):

*Hint: the slope is -1

OpenStudy (anonymous):

Now get one of the points (7,-6) or (6,-5). and plug it into the point slope formula. the point slope formula is y – y1 = m(x – x1). Replace the x1 and y1 with any coordinate point.

OpenStudy (vortish):

ok i see you subtracted both sets of cords to come up with -1 right

OpenStudy (anonymous):

yes. (y2-1)/(x2-x1)= slope.

OpenStudy (anonymous):

and replace the m with -1

OpenStudy (vortish):

so i come up with (y-y1)2x right

OpenStudy (anonymous):

Solve for y. Which point did you use? (7,-6) or (6,-5)?

OpenStudy (vortish):

ok im confused

OpenStudy (anonymous):

I'll use the point (7,-6) (y-(-6))=-1(x-7) Now solve for y (y+6)=-x+7 y+6=-x+7 y=-x+1

OpenStudy (vortish):

ok i get that so now what

OpenStudy (anonymous):

y=-x+1 is your answer.

OpenStudy (vortish):

ok and i see how you got to that so thanks

OpenStudy (vortish):

acording to the site it is wrong

OpenStudy (vortish):

they want you to write an equation

OpenStudy (anonymous):

There must be a mistake on your side. You can check by plugging any of the points into the equation.